OpenStudy (kathydimieri12):

If a sprinkler waters 1 over 8 of a lawn in 1 over 4 hour, how much time will it take to water the entire lawn? A.3/8 hour B.1/2 hour C.2 hours D.4 hours

OpenStudy (calculusxy):

\[\large \frac{ \frac{ 1 }{ 8 } }{ \frac{ 1 }{ 4 } } = \frac{ 1 }{ x }\] Notice how I created a proportion in here. In the first fraction, I have rate that we are implicitly given). With that rate, we can assume and figure out the amount of hours it will take to fill the entire lawn (whichI wrote down with \(1\)).

OpenStudy (calculusxy):

Now you are left to solve for \(\large x\).

OpenStudy (kathydimieri12):

1/2

OpenStudy (calculusxy):

Let's check: \[\frac{ 1 }{ 4 } = \frac{ 1 }{ 8 }x \rightarrow \frac{ 1 }{ 4 } \times \frac{ 8 }{ 1 } = \frac{ 8 }{ 4 } = 2\]

OpenStudy (calculusxy):

Do you agree?

OpenStudy (kathydimieri12):

ooo so two hours?

OpenStudy (calculusxy):

Yes! \(\large x = 2\)

OpenStudy (kathydimieri12):

thank you!

OpenStudy (calculusxy):

You're welcome !

OpenStudy (kathydimieri12):

i have one question ? what does it mean when it says "close question "

OpenStudy (calculusxy):

I think it means that there will only be two answers: yes or no.

OpenStudy (mrnood):

it means it will no longer appear onth elist of questions requiring help